OpenStudy (anonymous):

The function f(t) = 4t2 - 8t + 6 shows the height from the ground f(t), in meters, of a roller coaster car at different times t. Write f(t) in the vertex form a(x - h)2 + k, where a, h, and k are integers, and interpret the vertex of f(t).

OpenStudy (anonymous):

f(t) = 4(t - 1)2 + 3; the minimum height of the roller coaster is 3 meters from the ground f(t) = 4(t - 1)2 + 3; the minimum height of the roller coaster is 1 meter from the ground f(t) = 4(t - 1)2 + 2; the minimum height of the roller coaster is 2 meters from the ground f(t) = 4(t - 1)2 + 2; the minimum height of the roller coaster is 1 meter from the ground

OpenStudy (anonymous):

do you know how to write \[f(t)=4t^2-8t+6\] in vertex form?

OpenStudy (anonymous):

Um, that's what i need help with =)

OpenStudy (anonymous):

all your answer choices look the same except two have \(+2\) at the end and two have \(+3\)

OpenStudy (anonymous):

if you want to know which one is right, find \(f(1)\)

OpenStudy (anonymous):

Right =) It will be one of the +2 answers =)

OpenStudy (anonymous):

yes it will

OpenStudy (anonymous):

the minimum height is the output, it is \(f(1)=2\)

OpenStudy (anonymous):

Right that's what i got =)

OpenStudy (anonymous):

So it's C?

OpenStudy (anonymous):

as always, it is C

OpenStudy (anonymous):

=) Thanks so much, for your help =)

OpenStudy (anonymous):

yw